# Daco: Towards Application-Driven and Comprehensive Data Analysis via Code Generation

### v1

@@ Appendix B Implementation Details
For zero-shot API-based systems including ChatGPT and GPT-4, we evaluate two settings, directly reading the table content, and using code generation. For the former setting, we linearize the table content into text representation as model input. Due to token limit, we feed the first 20 rows as input, which covers the full content of 93% tables. For the code generation setting, we employ the pipeline described in Figure 5. When the generated code causes a syntax or runtime error, we re-sample the model until the generated code can be executed. We allow up to 5 resamplings for each turn. We use the gpt-3.5-turbo-16k-0613 API for ChatGPT and gpt-4-32k API for GPT-4. We limit the number of total coding turns maximally at 9. For annotation generation where GPT-4 self-correction is allowed, we limit the number of self-correction within 2 for each turn and 4 for the whole session.
@@
For finetuned models including Daco-RL and SFT, we use CodeGeeX2-6B (Zheng et al., 2023a) as the base model. We first train the SFT model using GPT-4 annotations, and then train our Daco-RL model on top of the SFT model. When training and , we initialize the model from the SFT model. When training our Daco-RL model, we initialize the value model from , and initialize the policy model from the SFT model. In inference, we use nucleus decoding with p = 0.9 and temperature = 1.0. Similarly, we allow up to 5 resamplings when the generated code causes an error. The SFT model is trained with 8 A100 GPU for about 4 hours. The Daco-RL model is trained with 8 A100 GPU for about 18 hours. Detailed hyper-parameters are in Table 7. The only hyper-parameter we tune is for Daco-RL. We experiment with 0.8, 0.9 and 1.0 and discover that 1.0 works the best.
@@
For experiments on external test sets, we directly evalute the trained SFT and Daco-RL model on InfiAgent-DA (Hu et al., 2024) and FeTaQA (Nan et al., 2022) test sets. For InfiAgent-DA, following the original paper, we add a reformatting step to reformat the generated data analysis report into the key-value format. We use ChatGPT to perform reformatting with a simplified prompt without in-context examples.
Appendix C Qualitative Examples
We show final answers generated by SFT and Daco-RL in Figure 6. Daco-RL better focuses on user query, while SFT tends to display generic statistics that are less relevant to user query.
@@ Appendix C Qualitative Examples
Appendix D GPT Prompts
Here we show the prompts we use for ChatGPT and GPT-4. Prompt for query generation is in Table 8. Prompt for helpfulness annotation collection is Table 9. Prompts for helpfulness and informativeness evaluation are Table 10 and 11.
@@ Appendix D GPT Prompts
I have a database of [database title]. As a [stakeholder role], I want to [describe intention]. Below are the intermediate steps of their analysis. Which analysis is more informative? The more informative analysis should produce execution results that stick relevant to my analysis goal and bring more insights to my analysis. Your response should be in the following format. Note: answer should be either Analysis-1 or Analysis-2 * Answer: answer * Reasoning: explain your reasoning here The reports are as follows: # Analysis-1 [analysis 1] # Analysis-2 [analysis 2] |

### v2

@@ Appendix D Implementation Details
For zero-shot API-based systems including ChatGPT and GPT-4, we evaluate two settings, directly reading the table content, and using code generation. For the former setting, we linearize the table content into text representation as model input. Due to token limit, we feed the first 20 rows as input, which covers the full content of 93% tables. For the code generation setting, we employ the pipeline described in Figure 2(b) in the main content. When the generated code causes a syntax or runtime error, we re-sample the model until the generated code can be executed. We allow up to 5 resamplings for each turn. We use the gpt-3.5-turbo-16k-0613 API for ChatGPT and gpt-4-32k API for GPT-4. We limit the number of total coding turns maximally at 9. For annotation generation where GPT-4 self-correction is allowed, we limit the number of self-correction within 2 for each turn and 4 for the whole session.
@@
For finetuned models including SFT, RLHF and fine-grained RLHF, we use CodeGeeX2-6B [44] as the base model. We first train the SFT model using GPT-4 annotations, and then train our RLHF models on top of the SFT model. When training and , we initialize the model from the SFT model. When training our fine-grained RLHF model, we initialize the value model from , and initialize the policy model from the SFT model. In inference, we use nucleus decoding with p = 0.9 and temperature = 1.0. Similarly, we allow up to 5 resamplings when the generated code causes an error. The SFT model is trained with 8 A100 GPU for about 4 hours. The RLHF models are trained with 8 A100 GPU for about 18 hours. Detailed hyper-parameters are in Table 6. The only hyper-parameter we tune is for fine-grained RLHF. We experiment with 0.8, 0.9 and 1.0 and discover that 1.0 works the best.
@@
Appendix E Details of LLM Evaluation
We adopt LLM for automatic evaluation due to the complexity of evaluating answer helpfulness. Prior work has shown that LLMs can reliably evaluate text generation quality based on task description and judging criteria [21]. However, it has been known that LLM evaluators favor their own generations [30]. To mitigate this potential bias, we evaluate model helpfulness with multiple LLM evaluators, including OpenAI’s gpt-3.5-turbo-0613 (now deprecated) and gpt-4o-mini-2024-07-18, Anthropic’s claude-3-5-sonnet-20240620, and meta-llama/Meta-Llama-3-8B-Instruct from Llama 3 family. The detailed evaluation results are in Table 7. We report the average scores from GPT-4o mini, Claude 3.5 Sonnet and Llama 3 8B as the final score. Additionally, we show that different LLM evaluators agree with each other, achieving a moderate to substantial inter-annotator agreement of 0.45 Cohen’s kappa and a very high Spearman correlation of 0.90 in model performance ranking. This further verifies the robustness of LLM-based evaluation.
| TestA | TestH | ||||||||
| Code gen | GPT-3.5 | GPT-4o | Claude 3.5 | Llama 3 | GPT-3.5 | GPT-4o | Claude 3.5 | Llama 3 | |
| TAPAS | ✗ | 25.00 | 17.61 | 17.08 | 22.89 | 24.50 | 15.00 | 15.00 | 19.50 |
| TAPEX | ✗ | 14.79 | 17.78 | 7.39 | 20.07 | 6.00 | 9.50 | 3.50 | 14.00 |
| ChatGPT | ✗ | 25.18 | 20.60 | 18.31 | 19.01 | 18.50 | 16.00 | 9.00 | 15.50 |
| GPT-4 | ✗ | 30.81 | 30.81 | 32.39 | 28.17 | 24.00 | 21.00 | 23.00 | 17.50 |
| ChatGPT | ✓ | 35.74 | 22.89 | 24.70 | 31.93 | 27.27 | 18.69 | 19.19 | 26.26 |
| GPT-4 | ✓ | 52.00 | 51.09 | 52.18 | 49.09 | 41.88 | 45.31 | 39.58 | 46.88 |
| SFT | ✗ | 21.51 | 19.72 | 12.68 | 24.47 | 9.50 | 11.50 | 5.00 | 17.50 |
| SFT | ✓ | 20.95 | 12.38 | 7.38 | 21.43 | 11.54 | 5.76 | 7.05 | 16.67 |
| RLHF | ✓ | 15.18 | 14.51 | 3.13 | 14.29 | 8.79 | 7.14 | 0.55 | 14.84 |
| FG-RLHF | ✓ | 28.54 | 20.71 | 9.51 | 28.05 | 21.05 | 13.82 | 7.24 | 16.45 |
Appendix F License Information
Based on Kaggle’s policy, it is allowed to use and redistribute datasets as long as adhering to each dataset’s specific license. All datasets utilized in this study are publicly accessible. A majority, 65%, are covered under various standardized licenses, all of which permit data redistribution for academic purposes. All these licenses allow data redistribution for academic purposes. The remaining 35% are licensed on an ad-hoc basis. We have manually reviewed their Kaggle descriptions to ensure there are no restrictions against using their data. However, since some datasets have limited description or licensing information, we will continue to monitor their status and will remove any dataset if issues arise. Detailed license information is listed in Table 8.
| License | Count |
| CC0-1.0 | 123 |
| Ad-hoc | 110 |
| DbCL-1.0 | 14 |
| Attribution 4.0 International (CC BY 4.0) | 14 |
| CC-BY-NC-SA-4.0 | 13 |
| CC-BY-SA-4.0 | 7 |
| ODbL-1.0 | 6 |
| GPL-2.0 | 4 |
| CC BY 4.0 | 4 |
| Community Data License Agreement - Permissive - Version 1.0 | 3 |
| Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) | 3 |
| MIT License | 2 |
| Open Government Licence v3.0 | 2 |
| ODC Public Domain Dedication and Licence (PDDL) | 2 |
| ODbL | 1 |
| UN Data License | 1 |
| ODC Attribution License (ODC-By) | 1 |
| Creative Commons Attribution 4.0 International License | 1 |
| Open Use of Data Agreement v1.0 | 1 |
| Attribution-NonCommercial-ShareAlike 3.0 IGO (CC BY-NC-SA 3.0 IGO) | 1 |
| Government Open Data License - India | 1 |
Appendix G Qualitative Examples
We show a few examples of input databases and their associated queries in Figure 8. Though the queries are automatically generated, they are of high quality, diverse, and relevant to the databases.
We show final answers generated by SFT and RLHF in Figure 9. RLHF better focuses on user query, while SFT tends to display generic statistics that are less relevant to user query.
@@ Appendix G Qualitative Examples
Appendix H Prompts
Prompt for query generation is in Table 9. Prompt for helpfulness annotation collection is Table 10. Prompt for helpfulness evaluation is Table 11.